OpenStudy (anonymous):

Find a real number,K , such that the line 20x + ky-12= 0 has y -intercept -19 .

OpenStudy (anonymous):

\[0=20x+ky-12\] sub in the yintercept (0,-19) \[0=20(0)+k(-19)-12\] \[0=-19k-12\] \[12=-19k\] \[k=\frac{ 12 }{ -19 }\]